OpenStudy (anonymous):

Solve for x: 8x + 10 - 5x = 15.

OpenStudy (anonymous):

8x=10-5x=15 combine x: 8x-5x = 3x then what do you do next to get x by itself

OpenStudy (anonymous):

3x +10 =15 -10 both sides 3x = 5 / 3 x = 5/3